# Remote Sensing Image Intelligent Interpretation with the Language-Centered Perspective: Principles, Methods and Challenges

### v1

@@ Abstract
The mainstream paradigm of remote sensing image interpretation has long been dominated by vision-centered models, which rely on visual features for semantic understanding. However, these models face inherent limitations in handling multi-modal reasoning, semantic abstraction, and interactive decision-making. While recent advances have introduced Large Language Models (LLMs) into remote sensing workflows, existing studies primarily focus on downstream applications, lacking a unified theoretical framework that explains the cognitive role of language. This review advocates a paradigm shift from vision-centered to language-centered remote sensing interpretation. Drawing inspiration from the Global Workspace Theory (GWT) of human cognition, We propose a language-centered framework for remote sensing interpretation that treats LLMs as the cognitive central hub integrating perceptual, task, knowledge and action spaces to enable unified understanding, reasoning, and decision-making. We first explore the potential of LLMs as the central cognitive component in remote sensing interpretation, and then summarize core technical challenges, including unified multimodal representation, knowledge association, and reasoning and decision-making. Furthermore, we construct a global workspace-driven interpretation mechanism and review how language-centered solutions address each challenge. Finally, we outline future research directions from four perspectives: adaptive alignment of multimodal data, task understanding under dynamic knowledge constraints, trustworthy reasoning, and autonomous interaction. This work aims to provide a conceptual foundation for the next generation of remote sensing interpretation systems and establish a roadmap toward cognition-driven intelligent geospatial analysis.
Index Terms:
Language-Centric Remote Sensing Interpretation, Global Workspace Theory, Multimodal Representation, Knowledge Association, Reasoning and Decision-Making.
@@ I Introduction
Language and vision, as two fundamental paradigms of human cognition, have maintained a dialectical relationship throughout the development of civilization. Martin Jay[1] reveals the profound impact of the privileging of vision on modern cognitive science, where the visual system, due to its intuitive spatial perception, has been granted epistemological primacy within the framework of empiricist cognition.In contrast, the centrality of language is rooted in the Platonic tradition of idealism and underwent a modern transformation through the works of Descartes, Kant, and ultimately Wittgenstein’s ”linguistic turn”, which emphasized the abstract capacity of language as the foundation of rational cognition. The interplay between these two cognitive paradigms is also reflected in the history of biological evolution: the emergence of visual organs triggered the Cambrian explosion of biodiversity, while the birth of language marked the beginning of the human intelligence revolution. Kant, in his famous dictum ”Thoughts without content are empty, intuitions without concepts are blind.”[2] revealed the complementarity between the two: Vision offers the richness of concrete perception, while language provides the stability of abstract representation. Together, they form a closed loop through which humans cognitively engage with the physical world. In the interdisciplinary field of remote sensing science and artificial intelligence, this duality of cognitive paradigms is mirrored in the evolutionary trajectory of image interpretation technologies. Current mainstream intelligent remote sensing interpretation systems follow a vision-centered paradigm, relying on neural networks to extract pixel-level features and achieving remarkable progress in single-task scenarios such as object detection[3] and land-cover classification[4].
However, this paradigm faces three fundamental limitations. First, multimodal tasks—such as cross-modal retrieval and reasoning-based decision-making—often rely on the stacking of separate models, lacking a truly unified representational space. The root cause lies in the vision-centered paradigm’s dependence on spatiotemporal alignment across different modalities. For example, X Sun et al.[5] trained a multimodal model using jointly aligned optical, SAR, and nDSM data to ensure that information from different modalities could align and fuse within the same spatiotemporal reference frame. However, in most multimodal task scenarios, inherent mismatches across modalities—such as differences in acquisition methods, temporal synchronization, and spatial resolution—make it extremely challenging to construct a shared, unified representational space. Second, the interpretation process exhibits a unidirectional open-loop characteristic, lacking a dynamic feedback mechanism with the physical world. Due to the difficulty of effectively verifying the accuracy of results in an end-to-end interpretation approach, the model cannot adaptively adjust based on real-world feedback during the inference process. This lack of closed-loop validation mechanisms leads to cumulative errors when the model faces complex and variable real-world scenarios, impacting its robustness and generalization ability. Third, visual representations are constrained by the ”seeing is believing” perceptual limitations, making it difficult to extract deeper semantic relationships beyond overt features. CREPE[6] research indicates that visual models struggle to capture the implicit semantic relationships (such as functional attributes and conceptual links) between targets during recognition, limiting their abstract reasoning abilities. This limitation creates inherent deficiencies in visual models when it comes to understanding complex scenes, reasoning about causal relationships, or performing high-level cognitive tasks, thereby restricting their knowledge expression and reasoning capabilities in multimodal fusion. At its core, these limitations stem from the neglect of the conceptualization and structuring capabilities inherent in the language system, causing the interpretation results to remain at the surface level of vision and failing to form a closed-loop cognitive chain of ”Perception-Cognition-Decision”.
Figure 1: Overview of the organizational structure of review
With the advancement of large language model (LLM) technologies, research[7, 8] has shown that integrating vision and language models can help establish conceptual relationships among visual features. This insight has spurred exploration into remote sensing multimodal large language models (MLLMs), such as RS-LLaVA[9] and EarthGPT[10], which aim to combine visual information with language models to enhance the unification of task interfaces. These models leverage vision-language alignment techniques, enabling remote sensing image interpretation systems to effectively integrate multi-source data and support query and explanation through natural language interaction—thereby bridging the semantic understanding limitations of traditional methods. In recent years, remote sensing image interpretation has been gradually evolving toward multimodal reasoning models, such as EarthMarker[11], which aim to endow interpretation systems with stronger reasoning and decision-making capabilities. Existing review articles[12, 13, 14] have summarized this trend, focusing primarily on the technological development of MLLMs and their performance across various remote sensing interpretation tasks, highlighting improvements in accuracy and applicability. However, most of these reviews remain at the level of task-oriented application summaries and lack an in-depth analysis of the interaction mechanisms among different modalities, particularly the characteristics of multimodal interaction centered on the language modality.
Therefore, this review aims to fill this gap by systematically exploring the potential of a language-centered paradigm in remote sensing image interpretation. Inspired by the Global Workspace Theory[15] (GWT) of human cognitive models, we propose a theoretical framework for a global workspace-driven, cyclically interactive interpretation system. This framework demonstrates how language can be used to organize the interactions among interpretation tasks, sensors, and models, thereby achieving a cognitive closed loop that overcomes the limitations of current interpretation systems. It identifies three major challenges faced by current technologies in constructing a language-centered intelligent interpretation framework: the bottleneck of unified multimodal representation, the difficulty of multi-source knowledge association, and the limited capabilities in reasoning and decision-making. This work provides a theoretical foundation for building language-centered remote sensing interpretation systems and seeks to chart a path for multimodal remote sensing intelligence that ”returns to the essence of cognition.” The main contributions of this paper are as follows:
@@
    Rethinking Remote Sensing Image Interpretation from a Cognitive Perspective: Starting from the human cognitive process, this work analyzes the cognitive limitations of current remote sensing image interpretation technologies and emphasizes a language-centered ”Perception–Cognition–Decision” paradigm, laying a theoretical foundation for future research.
@@
    • The lack of effective mechanisms for associating multi-source knowledge;
    • The limited reasoning and decision-making capabilities.
    Potential research directions to address these challenges are also discussed.
@@
    Outlook and Emerging Research Opportunities: Finally, this review summarizes and discusses the new opportunities brought by the paradigm shift toward language-centered interpretation frameworks in the field of remote sensing image understanding. It also proposes potential research directions, including novel areas not covered by previous surveys.
The overall structure of this review is illustrated in the Fig.1.
@@ II Language-Centered
Remote Sensing Intelligence Interpretation
II-A Global Workspace-Driven Interpretation Mechanism
Human cognition of the physical world is a cyclical and interactive process, with bottom-up (perception forming cognition) and top-down (cognition guiding perception) pathways being the two fundamental ways humans understand the world. These processes are unified through consciousness. How consciousness bridges the physical world and human cognition through these two pathways is explained by various theories. The GWT is one of the mainstream perspectives, proposed in 1998 by Baars[15]. It posits that there exists a ”Global Workspace” in the brain as a shared platform for gathering and interacting information, capable of integrating information from different sources to support complex cognitive functions. Within this framework, bottom-up sensory information and memories in the brain converge and associate in the global workspace, where reasoning and further action decisions are made, guided by human intention. Ultimately, the brain sends out action commands that control the body’s interaction with the external world, enabling a response to the complex environment. For example, when a person is driving a car to find a parking space, first, visual perception detects road signs, parking spaces, and other vehicles’ positions, auditory perception senses traffic noise, and tactile perception via the steering wheel provides feedback on the road surface. These sensory inputs are transmitted to the brain via the nervous system in a bottom-up manner and integrated in the global workspace. In the global workspace, reasoning driven by human intention (to find an available parking space) makes decisions based on existing knowledge (surrounding environment information, current vehicle position), determining the next actions (steering the wheel, pressing the brake/accelerator, etc.). This action leads to environmental changes, which in turn provide the brain with new sensory information, creating a continuous cycle of interaction.
As shown in Fig.2, in the human cognitive process, multi-source information from complex physical scenes is captured by sensory organs (such as the eyes, ears, skin, etc.) and transmitted to the human cognitive global workspace. The global workspace generates intervention intentions based on this sensory information and interacts with the physical world through effectors (such as hands and feet). In this interaction, the physical space and cognitive space influence each other, continually driving human thinking and learning processes. Throughout this process, humans continuously refine their skills, gradually improving their ability to handle complex tasks.
Similarly, remote sensing intelligence interpretation of the physical world is also a cyclical and interactive process. Temporal and spatial dynamics, along with the perception of complex physical scenes that are constantly changing, are captured by remote sensing observation systems and quantified into digital representations. After analyzing and modeling these data, the information from the physical scene can be transmitted to the global workspace of remote sensing intelligence interpretation. Under the decision support of the global workspace, the system issues control commands to the remote sensing observation devices, actively acquiring more sensory data based on the needs of the interpretation task, thereby continuously adjusting and optimizing the interpretation of the physical scene.
In general, the remote sensing observation system transforms physical information from the real world into digital representations through a bottom-up perception process, transmitting it to the interpretation space for processing. When these sensory inputs are insufficient to meet the demands of the interpretation task, the remote sensing intelligent interpretation model needs to adjust the observation strategy top-down based on the current interpretation requirements to further acquire sensory data. For example, in a ship target interpretation task, the remote sensing intelligent interpretation system captures images of ships on the sea surface through sensors mounted on drones and inputs the digitalized representations into the system. The intelligent interpretation system analyzes these data and, combined with task knowledge, captures key features (such as the hull number) for identifying the ship type. However, when the hull number cannot be clearly obtained due to occlusion from the viewpoint, the system directs the drone to interact with the physical world by moving, adjusting the focal length, or switching sensory modalities, performing targeted actions to capture key details of the hull number.
@@ II-A Global Workspace-Driven Interpretation Mechanism
The cyclic and interactive interpretation process includes a bottom-up remote sensing image acquisition procedure and a top-down reasoning and decision-making process guided by the interpretation task. This process can be driven by a Global Workspace and formally represented as Cyclic Interaction Interpretation:
@@
The entire process can be abstracted as shown in Fig.3, where:
  • $P$ representing the perception space, denotes information such as scene context, object characteristics, and spatiotemporal attributes. This space captures a detailed description of geo-objects within the observation environment, including their spectral signatures, geometric properties, texture features, and temporal dynamics.
  • $K$ represents the knowledge space, referring to supplementary information derived from remote sensing expert knowledge. This includes prior knowledge of object features, geographical distribution patterns, seasonal variation rules, and more. Such knowledge can be acquired through expert experience or accumulated historical data, and is used during the interpretation process to enhance the understanding and reasoning over observed information.
  • $T$ denotes the task space, representing the user’s interpretation requirements—i.e., the specific objectives the interpretation system is expected to accomplish. Tasks within the task space may range from simple ones (such as identifying the object category in an image) to more complex ones (such as recognizing the fine-grained category of a ship in the image when key visual features are missing).
  • AA represents the action space, referring to the set of executable actions available to the sensing system. These actions include adjustments to the sensor’s resolution, intrinsic parameters, sensing modalities, combinations of modalities, as well as changes in observation angles and orientations.
  • $G$ denotes the Global Workspace, which integrates information from $P$ (Perception Space), $K$ (Knowledge Space), and $T$ (Task Space), and outputs control instructions to $A$ (Action Space).
In the cyclic interaction interpretation process, the Global Workspace encodes and integrates the visual representations from the Perception Space, the remote sensing expert knowledge from the Knowledge Space, and the interpretation requirements from the Task Space into a unified representation. It then establishes knowledge associations between perceptual information and expert knowledge. Guided by the interpretation requirements, Global Workspace performs reasoning and decision-making based on the perceptual information and expert knowledge, thereby controlling subsequent interactive actions.
@@ II-B The Potential of Applying LLMs as the Global Workspace
The global workspace drives the cyclic interaction interpretation process through its capabilities in unified encoding, knowledge association, and reasoning-based decision-making. With the rapid development of LLMs, such as LLaMA[16], GPT[17] and DeepSeek[18], these models have been shown to effectively encode world knowledge and exhibit logical reasoning capabilities. Meanwhile, the integration of LLMs has also led to significant advancements in areas such as unified multimodal modeling, knowledge base construction, and embodied intelligence.
@@
Therefore, this paper argues that LLMs possess strong potential to serve as the global workspace within remote sensing intelligent interpretation systems, for the following reasons:
LLMs possess the ability to construct unified representations across multimodal data
@@
Different modalities essentially express the same entity from distinct perspectives. Based on this view, the Platonic Representation Hypothesis suggests that representations across modalities tend to converge consistently, and using language as the central convergence point facilitates cross-modal information association and reasoning.In the visual domain, pioneering vision-language models such as CLIP[7] and ALIGN[19] jointly train vision and language modalities by mapping both into a shared semantic space. This significantly enhances the model’s ability to understand deep semantic associations between visual elements, confirming that although image and text are distinct in modality, they can converge through shared semantic concepts.BLIP[20] further supports this by generating descriptive textual captions for visual inputs, thereby demonstrating that visual information can be effectively described through language and enabling the construction of a shared vision-language representation space.In the broader multimodal context, LanguageBind[21] employs contrastive learning to align a variety of modalities—including video, infrared, depth, and audio—with language, enabling their embedding into a unified language-centered space. Furthermore, AllSpark[22] successfully aligns representations from 13 different modalities into the language modality.These works collectively validate: (1).That heterogeneous modality data can be embedded into a unified space via the language modality; (2).That diverse modality representation spaces can be integrated through a language-centered framework.
LLMs possess the ability to associate perceptual information with existing knowledge
@@
Through training on large-scale corpora, LLMs encode the logical relationships within language, manifesting in their strong logical reasoning capabilities[23, 24]. This enables LLMs to first comprehend the perceptual information in multimodal data and then match it with previously learned knowledge, thus driving advancements across multiple fields. In terms of information understanding, LLMs can not only understand tasks in algebra, calculus, and geometry by recognizing objectives and variable conditions, thereby constructing appropriate problem-solving logic, but also handle geographic data mining and analysis tasks, gaining deep insights into task objectives such as spatial patterns, geographic entity relationships, and trend analysis.Regarding knowledge association, the knowledge linking ability of LLMs stems from the vast world knowledge, linguistic patterns, and conceptual associations accumulated during their training process. This knowledge serves as the model’s ”long-term memory”, allowing it to quickly integrate relevant information and provide explanations when faced with complex inputs. The PaLM model proposed by Chowdhery et al.[25] demonstrates that the large-scale parameters of LLMs are crucial for enhancing retrieval capabilities. This allows the model to associate learned knowledge with minimal prompting in complex contexts, enabling it to make accurate responses.
LLMs possess the ability to handle complex task reasoning and decision-making
@@
Reasoning is essentially the process of drawing conclusions from known information or premises through logical inference chains. LLMs achieve reasoning and decision-making for complex tasks by constructing inference chains and progressively deducing from the associated information.Research by Wei et al.[26] explored the role of chain-of-thought prompting in large language models, demonstrating that by clearly defining reasoning links, LLMs can not only better understand the task itself but also efficiently solve complex tasks that require multiple reasoning steps. Moreover, reasoning and decision-making for complex tasks often require LLMs to dynamically adjust their reasoning chains based on task objectives. Relevant studies have shown that LLMs can optimize their decision-making process by adjusting their reasoning chains.An example of this capability is the DeepSeek-R1[27] model which can adjust its reasoning chain and refine its reasoning strategy before responding. By incorporating feedback information, it can correct previous errors, thereby solving complex tasks. For instance, during the AIME 2024 test, DeepSeek-R1[27] successfully solved 79.8% of high-difficulty problems by reflecting on and adjusting its reasoning steps.

### v2

# Remote Sensing Image Intelligent Interpretation from the Language-Centered Perspective: Principles, Methods and Challenges

@@ Abstract
Remote sensing image interpretation has predominantly followed a visual-centered paradigm, relying on visual feature extraction to achieve semantic understanding. However, faced with increasingly complex interpretation demands, such methods have inherent limitations in cross-modal reasoning, abstract semantic representation, and decision-making. Although large language models (LLMs) have recently been applied to the remote sensing domain, existing research focuses mostly on performance improvements in specific downstream tasks and lacks a systematic theoretical explanation of the core role of language in the cognitive process. This article aims to reveal the paradigm shift in remote sensing interpretation from “visual-centere” to “language-centered” Inspired by global workspace theory (GWT) in cognitive science, we construct a language-centered remote sensing interpretation framework. This framework positions the LLM as a cognitive hub, coordinating the four spaces of perception, task, knowledge, and action, thereby aiming to achieve a unified loop of understanding, reasoning, and decision-making. We first analyze the potential of LLM as core cognitive components and then systematically review key challenges in realizing this paradigm: multimodal representation, knowledge association and reasoning, and decision-making and execution. To address these challenges, we summarize existing solutions, training datasets and evaluation benchmarks. Finally, we outline future research directions regarding the adaptive alignment of multimodal data, task understanding under dynamic knowledge constraints, trustworthy reasoning, and autonomous interaction. This review aims to provide a theoretical perspective for understanding the role of language in remote sensing cognition and to serve as a reference for building next-generation intelligent cognitive geospatial analysis systems.
@@ I Introduction
Language and vision, as two fundamental paradigms of human cognition, have maintained a dialectical relationship throughout the development of civilization. Martin Jay[61] reveals the profound impact of the privileging of vision on modern cognitive science, where the visual system, owing to its intuitive spatial perception, has been granted epistemological primacy within the framework of empiricist cognition. In contrast, the centrality of language is rooted in the Platonic tradition of idealism and underwent a modern transformation through the works of Descartes, Kant, and ultimately, Wittgenstein’s linguistic turn, which emphasized the abstract capacity of language as the foundation of rational cognition. The interplay between these two cognitive paradigms is also reflected in the history of biological evolution: the emergence of visual organs triggered the Cambrian explosion of biodiversity, while the birth of language marked the beginning of the human intelligence revolution. Kant, in his famous dictum “Thoughts without content are empty, intuitions without concepts are blind.”[130], revealed the complementarity between the two: Vision offers the richness of concrete perception, while language provides the stability of abstract representation. Together, they form a closed loop through which humans cognitively engage with the physical world. In the interdisciplinary field of remote sensing science and artificial intelligence, this duality of cognitive paradigms maps onto the evolutionary trajectory of image interpretation technologies.
Current mainstream remote sensing intelligent interpretation systems adhere to a visual-centered paradigm: they rely on vision models to extract pixel-level or region-level features from imagery and accomplish tasks such as scene classification, object recognition, and target detection through the mapping between features and semantic labels and have achieved significant success in these standardized scenarios[208, 83]. However, as remote sensing interpretation tasks continuously expand in both form and reasoning depth, the information obtainable solely from visual signals has become increasingly insufficient. The visual-centered interpretation paradigm, limited by explicit and local perceptual cues, struggles to perform complex reasoning based on deep semantic relationships, leaving them stretched when facing demands for high-order object relationship mining and dynamic semantic reasoning.
TABLE I: Comparison with Related Reviews
Reviews Year Key Focus
Li et al.[86] 2024 Applications of VLMs: Focuses on applications in specific tasks such as VQA and image captioning.
Huang et al.[56] 2025 Mar The transition from vision to multimodality: Provides a detailed discussion of the technological evolution from CNN/ViT-based vision models to MLLMs.
Xiao et al.[181] 2025 Jun Remote Sensing Foundation Models: Integration of VFMs, VLMs, and LLMs; highlights general-purpose applicability and zero-shot transferability.
Lu et al.[104] 2025 Sep Vision Foundation Models: Focuses on VFM (e.g., SAM, MAE) for pixel-level and region-level perception tasks in remote sensing.
Ours - Language-Centered Interpretation Paradigm: Incorporates global workspace theory (GWT), positioning language as the cognitive hub to shift from passive interpretation to active interaction and decision-making.
To surmount the bottlenecks of the aforementioned visual-centered paradigm in terms of semantic reasoning and interaction capabilities, intelligent remote sensing interpretation technology has undergone a series of significant evolutions and iterations. Visual foundation models (VFMs)[157, 112, 104] significantly enhanced general feature representation capabilities through large-scale pretraining; however, they remained essentially confined to visual pattern matching, lacking the explicit expression of abstract concepts. Subsequently, the introduction of vision-language models (VLMs)[86] overcame this limitation, preliminarily achieving cross-modal semantic association by establishing alignment relationships between visual features and linguistic semantics. With breakthroughs in large language model technology, multimodal large language models (MLLMs)[56] emerged. Works such as RS-LLaVA[13] and EarthGPT[199] deeply integrated visual encoders with LLMs, utilizing the language modality to complement high-level semantic information missing in visual cues and unifying the interaction interface for multiple tasks. Recent research has further pushed this trend to the stage of LLM-based agents[184, 140]. For instance, models such as EarthMarker[198] treat the LLM as a cognitive hub, endowing the interpretation system with capabilities for proactive planning, complex logical reasoning, and decision-making. This technological evolutionary trajectory from perception (VFM) to alignment (VLM) and then to understanding (MLLM) and cognition (agent) reveals that remote sensing interpretation is undergoing a fundamental paradigm shift: transitioning from a “visual-centered paradigm” that relies on explicit visual features to a “language-centered paradigm” centered on semantic understanding and logical reasoning.
Although several recent reviews have systematically summarized new breakthroughs in the field of remote sensing image interpretation (as shown in TableI), they focus primarily on the categorization of technical implementations and performance comparisons. For instance, Lu et al.[104] focus on the architecture and pretraining strategies of VFMs, emphasizing pixel-level feature representation capabilities; Li et al.[86] and Huang et al.[56] discuss in detail the application of VLMs in multimodal alignment and downstream tasks (such as VQA and image captioning). However, most of these works remain at the level of “instrumental rationality”, focusing on “how to train stronger models” or “how to fuse more data”, while rarely offering reflections at the level of cognitive essence: Why does language occupy a central position in the evolution of remote sensing interpretation? What kind of cognitive paradigm improvement does the shift from visual feature mapping to linguistic logical reasoning truly signify?
Figure 1: Overall structure
To bridge this gap, this paper aims to establish a language-centered cognitive theoretical framework. We introduce global workspace theory (GWT)[11] from cognitive science, redefining the LLM from a mere “interaction interface” into a “cognitive hub”. The overall structure of this review is illustrated in Fig.1. Its main contributions are as follows:
@@
    Rethinking Remote Sensing Image Interpretation from a Cognitive Perspective: Starting from the human cognitive process, this review analyzes the cognitive limitations of current remote sensing image interpretation technologies and emphasizes a language-centered “perception-cognition-decision” paradigm, thereby laying a theoretical foundation for future research.
@@
    • The lack of effective mechanisms for combining multisource knowledge and reasoning;
    • The limited decision-making and execution capabilities.
@@
    Outlook and Emerging Research Opportunities: Finally, this review summarizes and discusses the new opportunities brought by this language-centered paradigm.
@@ II Language-Centered
Remote Sensing Intelligent Interpretation
II-A What Is Language-Centered Interpretation and Why Does It Matter?
Language-centered remote sensing (RS) interpretation is a paradigm that employs the language modality as a unified semantic interface to organize perception, task representation, and knowledge reasoning. As a symbolic system, language represents concepts, attributes, and relationships through discrete symbols and combinatorial rules, enabling information to be represented and manipulated in a structured and compositional manner. In cognitive studies, language is regarded as the container of human thought[19], providing the essential mechanism for concept formation, relationship building, and high-level reasoning, which allows knowledge to be organized and controlled at an abstract level. In 1998, Bernard Baars proposed global workspace theory (GWT)[11]. This theory posits that the brain contains a “global workspace” that serves as a shared platform for information convergence and interaction. It integrates content from diverse sources, such as perception and memory, to support complex cognitive functions. Within this framework, as shown in Fig. 2, perceptual information from complex physical scenes and knowledge stored in memory intersect and correlate within the workspace, driving reasoning processes and generating actionable decisions under the guidance of intent. Subsequent actions affect the external world, and the resulting environmental changes provide feedback to the cognitive system through new perceptual inputs, thereby forming a continuously updating cycle of interactive engagement.
Inspired by this cognitive framework, intelligent remote sensing (RS) interpretation can be viewed as a continuous, cyclic interactive process. Complex and dynamic spatiotemporal scenes are first captured by RS observation systems and transformed into digitized perceptual information. Once parsed and modeled, this information enters the “global workspace” of the interpretation system to support scene understanding and task reasoning (as shown in Fig. 2). Guided by the decisions generated within the workspace, the system issues commands back to the observation equipment. Based on the requirements of the interpretation task, it proactively acquires supplementary information to continuously refine and optimize its understanding of the external world. In this context, RS observation represents a bottom-up perceptual process, whereas RS interpretation employs top-down strategy adjustments to secure necessary information; together, they form a continuous closed loop. Within this closed-loop framework, language serves as the pivotal mechanism for task expression, semantic organization, and reasoning guidance. This feature enables the system to transcend the limitations of purely visual representations, endowing it with enhanced task-expressive power, deeper reasoning capabilities, and a more robust basis for decision-making.
@@ II-A What Is Language-Centered Interpretation and Why Does It Matter?
This process can be driven by a global workspace and formally represented as cyclic interaction interpretation:
@@
The entire process is shown in Fig.3, where
  • $P$ represents the perception space and denotes information such as scene context, object characteristics, and spatiotemporal attributes. This space captures a detailed description of geo-objects within the observation environment, including their spectral signatures, geometric properties, texture features, and temporal dynamics.
  • $K$ represents the knowledge space, referring to supplementary information derived from remote sensing expert knowledge. This information includes prior knowledge of object features, geographical distribution patterns, seasonal variation rules, and other details. Such knowledge can be acquired through expert experience or accumulated historical data and is used during the interpretation process to enhance the understanding and reasoning of observed information.
  • $T$ denotes the task space, which represents the user’s interpretation requirements. Tasks within the task space may range from simple (such as identifying the object category in an image) to more complex (such as recognizing the fine-grained category of a ship in the image when key visual features are missing).
  • AA represents the action space, which refers to the set of executable actions available to the sensing system. These actions include adjustments to the sensor’s resolution, intrinsic parameters, sensing modalities, combinations of modalities, and changes in observation angles and orientations.
  • $G$ denotes the global workspace, which integrates information from $P$ (perception space), $K$ (knowledge space), and $T$ (task space) and outputs control instructions to $A$ (action space).
In the cyclic interaction interpretation process, the global workspace encodes and integrates the visual representations from $P$, the remote sensing expert knowledge from $K$, and the interpretation requirements from $T$ into a unified representation. It first establishes associations between perceptual information and expert knowledge and performs reasoning based on these connections. Guided by the interpretation requirements, the global workspace then makes decisions based on the reasoning outcomes and finally executes the determined interactive actions in $A$.
@@ II-B The Potential of Applying LLMs as the Global Workspace
With the rapid development of LLMs, such as LLaMA[161], GPT[17] and DeepSeek[14], these models have been shown to effectively encode world knowledge and exhibit logical reasoning capabilities. This paper argues that LLMs possess strong potential to serve as the global workspace within remote sensing intelligent interpretation systems for the following reasons:
LLMs can construct unified representations across multimodal data
@@
The Platonic Representation Hypothesis[57] posits that data from different modalities are essentially expressions of the same underlying reality across different dimensions, exhibiting a tendency toward consistent convergence. Utilizing the language modality as the unified convergence point facilitates cross-modal information association and reasoning. In terms of visual modality representation, pioneering vision-language models such as CLIP[134], ALIGN[62], and BLIP[81] map visual and linguistic modalities into a shared semantic space through contrastive learning. These works effectively enhance the model’s deep understanding of object relationships, confirming that visual and language modalities can achieve consistent convergence through shared semantics. With respect to multimodal representation, LanguageBind[206] employs contrastive learning to align various modalities–including video, infrared, depth maps, and audio–with the language modality, hence achieving multimodal embedding into the latent space of LLM. Furthermore, AllSpark[142] embeds 13 distinct modalities into a language-centered representation.
LLMs can associate perceptual information with existing knowledge
@@
Through training on large-scale corpora, LLMs encode the logical relationships within language, as manifested in their strong logical reasoning capabilities[7, 2]. This feature enables LLMs to first comprehend perceptual information in multimodal data and then match it with previously learned knowledge, thus driving advancements across multiple fields. In terms of information understanding, LLMs can not only understand tasks in algebra, calculus, and geometry by recognizing objectives and variable conditions, thereby constructing appropriate problem-solving logic but also handle geographic data mining and analysis tasks, gaining deep insights into task objectives such as spatial patterns, geographic entity relationships, and trend analysis. With respect to knowledge association, the knowledge linking ability of LLMs stems from the vast amount of world knowledge, linguistic patterns, and conceptual associations accumulated during their training process. This knowledge serves as the model’s “long-term memory”, allowing the model to quickly integrate relevant information and provide explanations when faced with complex inputs. The PaLM model proposed by Chowdhery et al.[28] demonstrates that the large-scale parameters of LLMs are crucial for enhancing retrieval capabilities. This characteristic allows the model to associate learned knowledge with minimal prompting in complex contexts, thus enabling it to make accurate responses.
LLMs can handle complex task reasoning and decision-making
@@
Reasoning is essentially the process of drawing conclusions from known information or premises through logical inference chains [141, 185]. LLMs achieve reasoning and decision-making for complex tasks by constructing inference chains and progressively deducing them from the associated information. Wei et al.[176] explored the role of chain-of-thought prompting in large language models and demonstrated that by clearly defining reasoning links, LLMs can not only better understand the task itself but also efficiently solve complex tasks that require multiple reasoning steps. Moreover, reasoning and decision-making for complex tasks often require LLMs to dynamically adjust their reasoning chains based on task objectives. Relevant studies have shown that LLMs can optimize their decision-making process by adjusting their reasoning chains. An example of this capability is the DeepSeek-R1[42] model, which can adjust its reasoning chain and refine its reasoning strategy before responding. By incorporating feedback information, it can correct previous errors, thereby solving complex tasks. For instance, during the AIME 2024 test, DeepSeek-R1[42] successfully solved 79.8% of high-difficulty problems by reflecting on and adjusting its reasoning steps.